# Topological Magnon-Plasmon Hybrids

### v3

@@ Introduction.
Figure 1: Sketch of topological chiral edge magnon-plasmons in a bilayer of a metal (top layer) and a magnetic skyrmion crystal (bottom layer). The magnons of the skyrmion crystal are coherently coupled to the plasmons by magnetic dipole coupling, causing a topologically non-trivial spectral anticrossing in the bulk, which supports the chiral hybrid excitations at the edges of the sample (indicated by the water-wave-like feature and the yellow wave packet with an arrow).
@@
The full bilinear part of $H^{(1)}$ in Eq. (1) reads $H^{(1)}=\frac{1}{2}\sum_{\boldsymbol{k}}\boldsymbol{\phi}^{\dagger}_{\boldsymbol{k}}H_{\boldsymbol{k}}\boldsymbol{\phi}_{\boldsymbol{k}}$ where $\boldsymbol{\phi}^{\dagger}_{\boldsymbol{k}}=(b^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}},b_{-\boldsymbol{k}},a_{-\boldsymbol{k}})$. We drop the anomalous coupling ($a^{\dagger}b^{\dagger}$ and $ab$) between magnons and plasmons, because their coupling strength is much smaller than the energy at which the bands cross. Thus, the dynamics of the coupled system is approximated by $H^{(1)}\approx\sum_{\boldsymbol{k}}\boldsymbol{\psi}^{\dagger}_{\boldsymbol{k}}\tilde{H}_{\boldsymbol{k}}\boldsymbol{\psi}_{\boldsymbol{k}}$, with $\boldsymbol{\psi}^{\dagger}_{\boldsymbol{k}}=(b^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}})$ and
@@
As an example, we choose $S=5/2$, $K=0.04\,\mathrm{meV}$, $J=1.15\,\mathrm{meV}$ and $N_{z}=10^{3}$ and plot the magnon-plasmon anticrossing in Fig. 2(a), with color indicating the magnon and plasmon character. The emerging Berry curvature gives rise to an anomalous velocity and causes an intrinsic thermal Hall effect, that is, a heat current density $q_{x}=-\kappa_{xy}\partial_{y}T$ directed along the $x$ direction as response to a temperature gradient $\partial_{y}T$ in the $y$ direction, with the thermal Hall conductivity $\kappa_{xy}$ given by [38, 39, 40]
@@
where $k_{B}$ is the Boltzmann constant, $T$ is temperature, and $A$ is the total area of the system. Furthermore, $c_{2}(x)=(1+x)(\ln{\frac{1+x}{x}})^{2}-(\ln{x})^{2}-2\text{Li}_{2}(-x)$, where $\text{Li}_{2}(x)$ is the dilogarithm function, and $\rho_{n,\boldsymbol{k}}=[\mathrm{e}^{E_{n,\boldsymbol{k}}/(k_{B}T)}-1]^{-1}$ is the Bose-Einstein function. As shown in Fig. 2(b), the thermal Hall conductivity gets thermally activated and reaches a plateau around $k_{\text{B}}T\sim\Delta$. With typical FM resonance frequencies in the range of some tens of GHz, temperatures of a few Kelvin are enough to saturate $\kappa_{xy}$.
@@
$H_{\text{magnet}}^{(2)}=\frac{J}{2}\sum_{\langle i,j\rangle}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}-K\sum_{i}\left(S_{i}^{z}\right)^{2}$ (8)
@@
is the magnon spin Berry curvature, where we defined the spin current $\boldsymbol{j}^{S}=(S_{z}\boldsymbol{v}+\boldsymbol{v}S_{z})/2$ with velocity $\boldsymbol{v}=\hbar^{-1}\partial\tilde{H}_{\boldsymbol{k}}/\partial\boldsymbol{k}$ and spin $S_{z}=\hbar\,\text{Diag}(1,-1,0)$. As shown in Fig. 3(b), we find $\alpha_{xy}$ is non-zero, which is consistent with time-reversal symmetry. We conclude that the coupling to plasmons can be used to engineer spin currents in simple antiferromagnets.
@@
In both the ferromagnet and the antiferromagnet, we have found a bulk anticrossing, which in the ferromagnet give rise to nonzero Chern numbers. However, we do not expect to see well-defined chiral edge states because the bare dispersions, $\omega_{k}$ and $\Omega_{k}$, increase monotonously, resulting in a surface spectrum that does not have a band gap to support edge states. Thus, we next explore magnetic systems that naturally feature magnetically active magnon bands with negative curvature in the long-wavelength limit: magnetic skyrmion crystals (SkX). In particular, we focus on counterclockwise rotation (CCW) mode [44] with a negative effective mass.
Topological spin textures have been found in a variety of van der Waals magnets, including metals (Fe3GeTe2 [45], WTe2/Fe3GeTe2 heterostructures [46], and Cr2Ge2Te6 [47]) as well as insulators (CrBr3 [48]). Furthermore, a large Dzyaloshinskii-Moriya interaction was predicted even in a monolayer of Janus van der Waals magnets [49, 50, 51]. Inspired by these experimental observations and computational predictions, we consider skyrmion-hosting van der Waals magnets stacked on graphene and a dielectric spacer. A minimal spin-lattice model is defined on a triangular lattice as
@@
The SkX becomes the ground state due to the competition between ferromagnetic exchange and Dzyaloshinskii-Moriya interactions at finite magnetic fields but can be stabilized as a metastable state without external fields [45, 52]. We obtain the classical spin configuration of SkXs by Monte Carlo annealing, which is relaxed via Landau-Lifshitz-Gilbert simulations. Using linear spin wave theory [53], we compute the magnon band structure of the zero-field SkX as shown in Fig. 4(a). We adjust the scaling of the energy and wave vector by assuming $J=12$ meV, $D=1.2$ meV, $K_{z}=0.03$ meV, and the size of magnetic unit cell L=100L=100~nm [35].
@@
Since the CCW mode gives rise to the rotation of the in-plane magnetization [54, 35], the magnon-plasmon coupling induces a winding in hybridized wave functions with a topological gap. In addition, the topological spin textures of SkXs support the nontrivial magnon band topology [55, 56, 57]. While the CCW mode is characterized by C=1C=-1, the magnon-plasmon coupling induces $C=+1$ below the hybridization gap. Thus, chiral edge modes of magnons and magnon-plasmon hybrids propagate in opposite directions, which might result in backscattering. However, a significant mismatch in energy and wave vector of these edge modes suppresses scattering among them [35]. Consequently, both edge modes should remain robust, although they are not topologically protected.
@@
Figure 4: Topological magnon-plasmon polariton in a heterostructure of SkXs and graphene. (a) Three-dimensional magnon band structures of the metastable zero-field SkX. The energy and wave vector are rescaled from the dimensionless spin-lattice model as detailed in the SM [35]. (b,c) Local density of states (LDOS) of (b) magnons and (c) plasmons at an edge of the semi-infinite lattice. We set the damping rate of plasmons and magnons at $\kappa_{\textrm{pl}}/\omega_{\textrm{pl}}=\kappa_{\textrm{mag}}/\omega_{\textrm{mag}}=10^{-4}$.
We construct a two-band effective model to investigate the topological magnon-plasmon polaritons,
$H_{\textrm{eff}}$ $=\begin{pmatrix}\frac{\hbar^{2}k^{2}}{2M_{\textrm{mag}}}+E_{\textrm{mag}}&G\mathrm{e}^{-\mathrm{i}\varphi_{\boldsymbol{k}}}\\ G\mathrm{e}^{\mathrm{i}\varphi_{\boldsymbol{k}}}&\frac{\hbar^{2}k^{2}}{2M_{\textrm{pl}}}+E_{\textrm{pl}}\end{pmatrix}$, (13)
where $M_{\textrm{mag/pl}}$ is the effective mass, $E_{\textrm{mag/pl}}$ is the constant energy shift, and $G/h=0.55$ GHz. Here, we fit the dispersion relations of magnons and plasmons by quadratic terms, which reproduces the group velocity and frequency at $k=k_{c}$ [35]. Since the velocity of magnons is much slower than that of plasmons, the above model is sufficient to study the topological nature of the magnon-plasmon hybrid. The tight-binding model is obtained by introducing a cutoff in the wave vector.
We employ the renormalization technique to study the local density of states (LDOS) in the semi-infinite geometry [35, 58, 59]. Figure 4 (b) and (c) show the LDOS of magnons and plasmons at an edge of the semi-infinite lattice, respectively. Both plots show the chiral edge state that connects the lower magnon branch with the upper plasmon branch. We note that the LDOS of magnons is much larger than plasmons due to the difference in their effective masses. Also, the bulk plasmon spectrum exists in a narrow range of wave vectors.
@@
We also investigate the impact of larger magnon damping rates $\kappa_{\textrm{mag}}$ while keeping $\kappa_{\textrm{pl}}=10^{-4}\omega_{\textrm{pl}}$ [35]. As the damping rate increases, it becomes harder to identify the chiral edge state due to the broadening of the bulk spectrum. We find that the chiral edge state is robust up to $\kappa_{\textrm{mag}}=10^{-3}\omega_{\textrm{mag}}\approx 38$ MHz. The damping rate of plasmons was recently measured to be $3.1\times 10^{-4}\omega_{\textrm{pl}}$ in the infrared light [60]. In addition, the damping rate of magnons in magnetic insulators is approximated as $\kappa_{\textrm{mag}}=\alpha\omega$, where the lowest value of the Gilbert damping $\alpha$ is known as $\alpha=3\times 10^{-5}$ at room temperature in yttrium iron garnet [61]. Thus, sufficiently small magnon and plasmon damping rates are attainable at cryogenic conditions.
@@
The hybrid nature of the chiral edge state is advantageous for experimental observations. Even though observing the chiral edge state has been challenging in topological magnonics, they have been successfully observed in topological plasmonics and photonics [62, 63]. For example, the plasmonic chiral edge state can be excited by gold antennas [64] and detected by the near-field microscropy [62, 60] and the transmission spectrum [65]. In addition, magnonic components can be measured by the NV center magnetometry [66, 67] and the near-field Brillouin light scattering [68].
@@
We have identified topologically nontrivial magnon-plasmon hybrids and chiral magnon-plasmon edge states in effectively two-dimensional stacks of van der Waals layers. The resulting nontrivial band geometry was shown to lead to transverse heat and spin transport and to chiral edge magnon-plasmon hybrids. Our effective theory and, in particular, the lattice regularization required to calculate the edge spectrum in Fig. 4(c,d), rely on the quasiparticle approximation and neglect the Landau damping in the electron-hole continuum at larger momenta. For open boundaries, the magnon-plasmon gap is filled with a continuum of states. We expect that the hybridization of single-particle states with the continuum will lead to breakdown effects similar to those discussed in Ref. 69 for interacting topological magnon insulators and in Ref. 70 for a generic bosonic toy model. However, as long as the energy scale of these effects is smaller than the topological bulk gap, our effective treatment captures the relevant single-particle physics, and an effective non-Hermitian description as in Ref. 71 could be pursued. To avoid the continuum altogether, one could use intrinsically undamped plasmon modes in narrow electron bands, e.g., in Moire graphene bilayers [72].

### v4

@@ Introduction.
Figure 1: Sketch of topological chiral edge magnon-plasmons in a bilayer of a metal (top layer) and a magnetic system (bottom layer). The magnons are coherently coupled to the plasmons by magnetic dipole coupling, causing a topologically non-trivial spectral anticrossing in the bulk, which supports the chiral hybrid excitations at the edges of the sample (indicated by the water-wave-like feature and the yellow wave packet with an arrow).
@@
The full bilinear part of $H^{(1)}$ in Eq. (1) reads $H^{(1)}=\frac{1}{2}\sum_{\boldsymbol{k}}\boldsymbol{\phi}^{\dagger}_{\boldsymbol{k}}H_{\boldsymbol{k}}\boldsymbol{\phi}_{\boldsymbol{k}}$ where $\boldsymbol{\phi}^{\dagger}_{\boldsymbol{k}}=(b^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}},b_{-\boldsymbol{k}},a_{-\boldsymbol{k}})$. We drop the anomalous couplings ($a^{\dagger}b^{\dagger}$ and $ab$) between magnons and plasmons, because their coupling strength is much smaller than the energy at which the bands cross, and thus they give negligible corrections to the hybridization. Thus, the dynamics of the coupled system is approximated by $H^{(1)}\approx\sum_{\boldsymbol{k}}\boldsymbol{\psi}^{\dagger}_{\boldsymbol{k}}\tilde{H}_{\boldsymbol{k}}\boldsymbol{\psi}_{\boldsymbol{k}}$, with $\boldsymbol{\psi}^{\dagger}_{\boldsymbol{k}}=(b^{\dagger}_{\boldsymbol{k}},a^{\dagger}_{\boldsymbol{k}})$ and
@@
As an example, we choose $S=5/2$, $K=0.04\,\mathrm{meV}$, $J=1.15\,\mathrm{meV}$ and $N_{z}=10^{3}$ and plot the magnon-plasmon anticrossing in Fig. 2(a), with color indicating the magnon and plasmon character. The emerging Berry curvature gives rise to an anomalous velocity and causes an intrinsic thermal Hall effect, that is, a heat current density $q_{x}=-\kappa_{xy}\partial_{y}T$ directed along the $x$ direction as response to a temperature gradient $\partial_{y}T$ in the $y$ direction, with the thermal Hall conductivity $\kappa_{xy}$ given by [38, 39, 40, 41]
@@
where $k_{B}$ is the Boltzmann constant, $T$ is temperature, and $A$ is the total area of the system. Furthermore, $c_{2}(x)=(1+x)(\ln{\frac{1+x}{x}})^{2}-(\ln{x})^{2}-2\text{Li}_{2}(-x)$, where $\text{Li}_{2}(x)$ is the dilogarithm function, and $\rho_{n,\boldsymbol{k}}=[\mathrm{e}^{E_{n,\boldsymbol{k}}/(k_{B}T)}-1]^{-1}$ is the Bose-Einstein function. As shown in Fig. 2(b), the thermal Hall conductivity gets thermally activated and reaches a plateau around $k_{\text{B}}T\sim\Delta$. With typical FM resonance frequencies in the range of some tens of GHz, temperatures of a few Kelvin are enough to saturate $\kappa_{xy}$. Here, we have dropped the anomalous couplings in the calculation of the thermal Hall conductivity, but we show in the SM [35] that including the anomalous couplings does not affect qualitatively the thermal Hall conductivity.
@@
$H_{\text{magnet}}^{(2)}=\frac{J}{2}\sum_{\langle i,j\rangle}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}-K\sum_{i}\left(S_{i}^{z}\right)^{2},$ (8)
@@
is the magnon spin Berry curvature, where we defined the spin current $\boldsymbol{j}^{S}=(S_{z}\boldsymbol{v}+\boldsymbol{v}S_{z})/2$ with velocity $\boldsymbol{v}=\hbar^{-1}\partial\tilde{H}_{\boldsymbol{k}}/\partial\boldsymbol{k}$ and spin $S_{z}=\hbar\,\text{Diag}(1,-1,0)$. As shown in Fig. 3(b), we find $\alpha_{xy}$ is non-zero, which is consistent with time-reversal symmetry. We conclude that the coupling to plasmons can be used to engineer spin currents in simple antiferromagnets. Including the anomalous couplings does not qualitatively affect the magnon spin Nernst conductivity at low temperatures, as shown in the SM [35].
@@
In both the ferromagnet and the antiferromagnet, we have found a bulk anticrossing, which in the ferromagnet gives rise to nonzero Chern numbers. However, we do not expect to see well-defined chiral edge states because the bare dispersions, $\omega_{k}$ and $\Omega_{k}$, increase monotonously, resulting in a surface spectrum that does not have a band gap to support edge states. Thus, we next explore magnetic systems that naturally feature magnetically active magnon bands with negative curvature in the long-wavelength limit: magnetic skyrmion crystals (SkX). In particular, we focus on counterclockwise rotation (CCW) mode [45] with a negative effective mass.
Topological spin textures have been found in a variety of van der Waals magnets, including metals (Fe3GeTe2 [46], WTe2/Fe3GeTe2 heterostructures [47], and Cr2Ge2Te6 [48]) as well as insulators (CrBr3 [49]). Furthermore, a large Dzyaloshinskii-Moriya interaction was predicted even in a monolayer of Janus van der Waals magnets [50, 51, 52]. Inspired by these experimental observations and computational predictions, we consider skyrmion-hosting van der Waals magnets stacked on graphene and a dielectric spacer. A minimal spin-lattice model is defined on a square lattice as
@@
The SkX becomes the ground state due to the competition between ferromagnetic exchange and Dzyaloshinskii-Moriya interactions at finite magnetic fields, but can be stabilized as a metastable state without external fields [46, 53]. We obtain the classical spin configuration of SkXs by Monte Carlo annealing, which is relaxed via Landau-Lifshitz-Gilbert simulations. Using linear spin wave theory [54], we compute the magnon band structure of the zero-field SkX as shown in Fig. 4(a). We adjust the scaling of the energy and wave vector by assuming $J=10$ meV, $D=1$ meV, $K_{z}=0.025$ meV, and the size of magnetic unit cell L=100L=100~nm [35].
@@
Since the CCW mode gives rise to the rotation of the in-plane magnetization [55, 35], the magnon-plasmon coupling induces a winding in hybridized wave functions with a topological gap. In addition, the topological spin textures of SkXs support the nontrivial magnon band topology [56, 57, 58]. While the CCW mode is characterized by C=1C=-1, the magnon-plasmon coupling induces $C=+1$ below the hybridization gap. Thus, chiral edge modes of magnons and magnon-plasmon hybrids propagate in opposite directions, which might result in backscattering. However, a significant mismatch in energy and wave vector of these edge modes suppresses scattering among them. Consequently, both edge modes should remain robust, although they are not topologically protected.
We construct the tight-binding model of magnon-plasmon hybrids in SkXs by fitting the dispersion of plasmons with quadratic terms [35]. Employing the renormalization technique for the semi-infinite geometry [59, 60], we compute the local density of states (LDOS) of magnon-plasmon hybrids at an edge of the semi-infinite skyrmion lattice as shown in Fig. 4(b) and (c), respectively. The damping rate is set at $\kappa_{\textrm{mag}}=\kappa_{\textrm{pl}}=4\textrm{ MHz}\approx 10^{-4}\omega_{\textrm{res}}$ with the resonance frequency $\omega_{\textrm{res}}=39$ GHz. Since the coupling strength between the CCW and plasmon modes is estimated as $G\sim 10^{-2}\omega_{\textrm{res}}$ [35], the strong coupling regime is realized in this setup. Figure 4(b) shows the LDOS over the one-dimensional first Brillouin zone of semi-infinite SkXs. The magnonic chiral edge mode is found inside the topological gap of the CCW mode, while it does not exhibit a clear signature of the anticrossing with plasmons. With a high resolution in energy and wave vector in Fig. 4(c), we obtain the magnon-plasmon edge state that connects the lower magnon branch with the upper plasmon branch. The LDOS of plasmon bands is not visible in Fig. 4(c) due to the significant difference in their effective masses. The group velocity of the magnon-plasmon edge mode is opposite to the magnonic edge mode above the CCW mode, which is consistent with the Chern number calculation.
@@
Figure 4: Topological magnon-plasmon polariton in a heterostructure of SkXs and graphene. (a) Three-dimensional magnon band structures of the metastable zero-field SkX. The energy and wave vector are rescaled from the dimensionless spin-lattice model as detailed in the SM [35]. (b,c) Local density of states (LDOS) of magnon-plasmon hybrids at an edge of the semi-infinite skyrmion lattice.
@@
We also investigate the effect of damping in the magnon-plasmon edge state [35]. As the damping rate of magnon modes increases, it becomes harder to identify the magnon-plasmon edge state due to the broadening of the bulk spectrum. We find that the chiral edge state is robust up to $\kappa_{\textrm{mag}}=10^{-3}\omega_{\textrm{mag}}$. In contrast, the magnon-plasmon edge state is less sensitive to the damping rate of plasmons, indicating the dominant magnonic contributions in the LDOS. The damping rate of plasmons was recently measured to be $3.1\times 10^{-4}\omega_{\textrm{pl}}$ in the infrared light [61]. In addition, the damping rate of magnons in magnetic insulators is approximated as $\kappa_{\textrm{mag}}=\alpha\omega_{\textrm{mag}}$, where the lowest value of the Gilbert damping $\alpha$ is known as $\alpha=3\times 10^{-5}$ at room temperature in yttrium iron garnet [62]. Thus, sufficiently small magnon and plasmon damping rates are attainable at cryogenic conditions.
@@
The hybrid nature of the chiral edge state is advantageous for experimental observations. Even though observing the chiral edge state has been challenging in topological magnonics, they have been successfully observed in topological plasmonics and photonics [63, 64]. For example, the plasmonic chiral edge state can be excited by gold antennas [65] and detected by the near-field microscropy [63, 61] and the transmission spectrum [66]. In addition, magnonic components can be measured by the NV center magnetometry [67, 68] and near-field Brillouin light scattering [69].
@@
We have identified topologically nontrivial magnon-plasmon hybrids and chiral magnon-plasmon edge states in effectively two-dimensional stacks of van der Waals layers. The resulting nontrivial band geometry was shown to lead to transverse heat and spin transport and to chiral edge magnon-plasmon hybrids. Our effective theory and, in particular, the lattice regularization required to calculate the edge spectrum in Fig. 4(b,c), rely on the quasiparticle approximation and neglect the Landau damping in the electron-hole continuum at larger momenta. For open boundaries, the magnon-plasmon gap is filled with a continuum of states. We expect that the hybridization of single-particle states with the continuum will lead to breakdown effects similar to those discussed in Ref. 70 for interacting topological magnon insulators and in Ref. 71 for a generic bosonic toy model. However, as long as the energy scale of these effects is smaller than the topological bulk gap, our effective treatment captures the relevant single-particle physics, and an effective non-Hermitian description as in Ref. 72 could be pursued. To avoid the continuum altogether, one could use intrinsically undamped plasmon modes in narrow electron bands, e.g., in Moire graphene bilayers [73].